Macrosoft, a profit maximizing software provider, has a fully enforceable patent on word processing software. They can produce software with no fixed costs and with a constant MC of $20 / software package. The market for word processing software is characterized by the demand curve:

QD = 60 - P

Macrosoft thus faces the following marginal revenue schedule: MR = 60 – 2Q

b. The newly elected mayor convenes his Council of Economic Advisors and hears the following testimony:

Professor A advises that if Macrosoft keeps its patent, because of its market power, it will be able to innovate in such a way that if it invests a fixed amount of $50, it will cut MC (marginal cost) in half at all levels of output.

If Professor A is correct, which of the following would be Macrosoft’s new TC (total cost) function? ?

a. TC = 50 + Q?

b. TC = 50*Q?

c. TC = 50 + 20*Q

d. TC = 50 + 10*Q ?

If Professor A is correct, Macrosoft would now maximize profit by producing _________ software packages and selling each for __________.

50 ; $10 / package?

25 ; $10 / package?

50 ; $35 / package?

25 ; $35 / package?

As a result of this innovation, Macrosoft’s profit would _________ to __________ from its level in (a).

Rise ; $875 ?

Rise ; $575 ?

Fall ; $250 ?

Fall ; Zero ?

iv. Total social surplus as a result of this innovation will be __________.

$887.50 ?

$937.50 ?

$800 ?

$650 ?
Answer #1

With no fixed cost, initially the total cost function was TC = 20Q. When it bears a fixed cost of $50, total cost function becomes TC = 50 + 20/Q/2 or TC = 50 + 10Q. Option D is correct

MR = 60 - 2Q and MC = 10 This implies MR = MC or 60 - 2Q = 10 or Q = 50/2 = 25 units with a price of P = 60 - 25 = $35. Option D is correct

Its profit will now be (35*25) - (50 + 10*25) = 575. Option C is correct

Total surplus = CS + PS = 0.5*(60 - 35)*25 + (35 - 10)*25 = 937.50. Option B is correct.